OpenStudy (ashking1):

Find the axis of symmetry of the graph of the function. f(x) = 2x2 - 4x

Nnesha (nnesha):

\[\huge\rm Ax^2+Bx+C\] quadratic equation you can use formula -b/2a to find axis of symmetry

OpenStudy (anonymous):

@AbbyWuvsChu talk to me

Nnesha (nnesha):

which is also the x-coordinate of the vertex a=leading coefficient b=coefficient of x term

OpenStudy (ashking1):

-4/4=-1

Nnesha (nnesha):

the formula is \[\rm \frac{ -b }{ 2a }\] and the b value is also negative

Nnesha (nnesha):

so it should be \[\frac{ -(-4) }{ 4 }\]

OpenStudy (ashking1):

1 oh ok

Nnesha (nnesha):

yes that's correct